# Isotope shift factors with quantum electrodynamics effects for many-electron systems: A study of the nuclear charge radius of 26mAl

### v1

@@ I Introduction
Until recently Plattner et al. (2023), the value of the ms charge radius for 26mAl was known only from extrapolations Towner and Hardy (2002, 2008). The study Plattner et al. (2023) resulted in a significant reexamination of the nuclear charge radius of 26mAl. However, the extraction of ms nuclear charge radii from these measurements requires input from atomic theory  Heylen et al. (2021); Plattner et al. (2023). The isotope shift between 26mAl and 27Al was measured using collinear laser spectroscopy of the $3s^{2}3p~{}^{2}P_{3/2}\to 3s^{2}4s~{}^{2}S_{1/2}$ transition Plattner et al. (2023), and the uncertainty in the extracted ms charge radii difference between 26mAl and 27Al was strongly dominated by the atomic theory calculations.
In the present paper, we undertake the accurate determination of atomic factors, which allows for a significant reduction in the theoretical and, hence, total uncertainty of the ms charge radius of 26mAl as well as 28Al, 29Al, 30Al, 31Al and 32Al isotopes. To achieve this, we perform high-precision relativistic electronic structure calculations within the coupled cluster theory with inclusion of up to quadruple excitation amplitudes. Furthermore, we develop an effective approach to incorporate leading-order quantum electrodynamics (QED) effects for these problems with precision comparable to that previously available only for similar transitions in highly charged ions, treated within the rigorous QED approaches.
@@ II.2 QED contribution to the field shift
Here $\theta(R_{\rm cut}-|\mathbf{r}|)$ is the Heaviside step function, $\eta_{kljm}(\mathbf{r})$ are H-like functions (with $n\leq 5$), and $R_{\rm cut}$ is a small radius, which can be varied slightly to study the stability of the results obtained. $X_{kljm,k^{\prime}ljm}$ in Eq. (10) are matrix elements of the SE operator over the $h_{kljm}$ functions and, for the reasons mentioned above, to good accuracy they are the linear combinations of the corresponding matrix elements over the H-like functions. Note, that the SE operator is diagonal in $ljm$. The expressions for both diagonal and off-diagonal matrix elements of SE operator can be derived, e.g., within the two-time Green’s function method Shabaev (2002):
$\langle\eta_{k}|\frac{1}{2}(\Sigma^{\rm SE}(\epsilon_{k})+\Sigma^{\rm SE}(\epsilon_{l}))|\eta_{l}\rangle,$ (12)
where $\Sigma^{\rm SE}(\epsilon_{k})$ is the renormalized SE operator. The model SE operator introduced in Ref. Shabaev et al. (2013) or its variant proposed in Ref. Skripnikov (2021) can be used to calculate the SE contribution to the total electronic energy in various systems Shabaev et al. (2013); Tupitsyn et al. (2016); Pašteka et al. (2017); Yerokhin et al. ; Machado et al. (2018); Si et al. (2018); Müller et al. (2018); Kaygorodov et al. (2019); Zaytsev et al. (2019); Shabaev et al. (2020); Skripnikov (2021); Kaygorodov et al. (2021); Savelyev et al. (2022); Kaygorodov et al. (2022); Savelyev et al. (2023); Skripnikov et al. (2021); Athanasakis-Kaklamanakis et al. (2023); Skripnikov and Barzakh (2024). However, they cannot be directly applied to evaluate the SE contribution to the FS constant (7).
@@
where, for brevity, we replaced the set of indices $kljm$ by one index $p$ (or $q$) for elements of the matrix $||X||$ in Eq. (10). We also explicitly indicated the dependence of these matrix elements on the ms nuclear charge radius, and $D_{p,q}^{\langle r^{2}\rangle}$ are elements of the one-particle density matrix in the same basis of functions $h_{i}$, which corresponds to the electronic wave function $\Psi^{\langle r^{2}\rangle}$. Now we have:
@@
To calculate SE contribution due to the first term of Eq. (15), we have to work out a procedure for computing the derivative of the SE matrix elements with respect to the nuclear ms charge radius. In Refs. Milstein et al. (2003, 2004); Yerokhin (2011); Jentschura (2003); Pachucki (1993); Milstein et al. (2002); Eides and Grotch (1997) the nuclear size dependence of the diagonal matrix elements of the SE operator over the functions of H-like ions has been explored and parameterized as follows:
@@
$=\frac{\alpha}{\pi}\Delta E_{\rm NS}(n\frac{1}{2}l,R)G_{\rm NSE}(njl,R),$ (16)
where we have omitted the total angular momentum projection in the index of the function $\eta$, $R=\sqrt{\frac{5}{3}}\sqrt{\langle r^{2}\rangle}$, $\Delta E_{\rm NS}$ is the nuclear size (NS) correction to the Dirac energy for the considered state $\eta_{nlj}$, and $G_{\rm NSE}$ is a slowly varying function of $Z$ and $R$ Yerokhin (2011). Let us consider the derivative of the matrix element (II.2):
$\frac{d\Delta E^{\rm SE}(nlj,R)}{d\langle r^{2}\rangle}\bigg{|}_{R_{0}}=\Delta E^{\rm SE}(nlj,R_{0})\left[\frac{\frac{d\Delta E_{\rm NS}(n\frac{1}{2}l,R)}{d\langle r^{2}\rangle}\bigg{|}_{R_{0}}}{\Delta E_{\rm NS}(nlj,R_{0})}+\frac{\frac{dG_{\rm NSE}(njl,R)}{d\langle r^{2}\rangle}\bigg{|}_{R_{0}}}{G_{\rm NSE}(nlj,R_{0})}\right],$ (17)
@@
where $\gamma=\sqrt{1-(\alpha Z)^{2}}$. Our numerical calculations within the hfd code Tupitsyn et al. (2002); Tupitsyn (2003) confirm this analytical result with high accuracy. The second term in the square brackets in Eq. (17) is independent of $n$ for $ns$ and $np_{3/2}$ states according to the explicit expression for $G_{\rm NSE}(njl,R)$ given in Ref. Yerokhin (2011). A certain dependence on $n$ does exist for $np_{1/2}$ states due to the dependence of the parameter $a_{21}=-2(n^{2}-1)/n^{2}$, which enters into the expression for $G_{\rm NSE}(njl,R)$ (see Eq. (10) of Ref. Yerokhin (2011)). To explore this dependence, we set $\sqrt{\langle r^{2}\rangle}\approx(2Z)^{1/3}+0.57$ fm approximately following the empirical dependence of $\sqrt{\langle r^{2}\rangle}$ on the nucleon number  Johnson and Soff (1985). We found that non-negligible dependence occurs in the region of $Z=36-40$. If we are outside this range, one can write:
@@ IV Results and discussion
Total $76.81(12)$ $76.80(6)$ $70.11(13)$ $70.10(8)$
Refs. Filippin et al. (2016); Heylen et al. (2021) $76.45(1)$.95)a $76.20(2)$.20)
@@
It is evident that the small uncertainties in the non-QED calculations of the atomic IS factors necessitate the consideration of the QED effects. These effects are typically overlooked in theoretical IS studies of many-electron neutral atoms due to uncertainties arising from the incomplete treatment of the correlation effects. Of particular interest is the SE QED contribution $F^{\rm SE}$ to the FS constant, $F^{\rm SE}$ which was found to be larger than the VP contribution (compare lines “QED-SE” and “QED-VP” of Table 5). In Section II.2, we proposed an effective method for performing such calculations for many-electron systems, which allows the use of the relativistic coupled cluster methods. The challenge arises from the lack of exact knowledge of the contribution of the finite nuclear size effect $\Delta E^{\rm SE}(nlj,R_{0})$ in Eq. (19). While the literature provides calculations for $Z=10$, 15, 20, etc. Shabaev et al. (2013), it does not cover $Z=13$. In the section II.2 we mentioned that the interpolation procedure can be used. The good agreement between the proposed technique and the rigorous QED treatment, as shown in Table 1 for Li-like ions, suggests that interpolation is viable. However, since Al is a light atom with $Z=13$, the interpolation scheme for it is based on only one point with a smaller $Z$ ($Z=10$). To verify the reliability of our approach, three calculations of the field shift constant were performed for the $3s^{2}3p~{}^{2}P_{1/2}\to 3s^{2}4s~{}^{2}S_{1/2}$ transition in Al: (i) employing Eq. (15) and the matrix elements $X^{\rm NS}_{kljm,k^{\prime}ljm}$ interpolated using the data from Ref. Shabaev et al. (2013), along with the analytical expressions for the derivative coefficients $M(lj,R_{0})$ obtained with the use of Eq. (17); (ii) using Eq. (15) with the directly calculated finite nuclear size contributions to the SE matrix elements and numerically computed $dX_{p,q}^{\langle dr^{2}\rangle}/d\langle r^{2}\rangle$ matrix elements within the ab initio QED methods Yerokhin and Shabaev (1999); Malyshev et al. (2022); Sapirstein and Cheng (2023); (iii) directly applying Eq. (7). In the latter case, two sets of the SE matrix elements over the H-like functions were specially calculated for Al for two different nuclear ms charge radii. Consequently, $F^{\rm SE}$ was calculated directly as a numerical derivative. The approach (iii) is the most accurate, while the approach (i) is the most approximate but relies solely on the QED data published in the seminal paper Shabaev et al. (2013) and does not require any new QED calculations. This significantly simplifies the problem of calculating the QED contribution, since an accurate evaluation of the SE matrix elements is not a trivial issue.

### v2

# Isotope-shift factors with quantum electrodynamics effects for many-electron systems: A study of the nuclear charge radius of 26mAl

@@ I Introduction
Nuclear ms charge radii can be determined through measurements of the spectral line isotope shifts (IS) between different isotopes of a given atom Yang et al. (2022). The corresponding technique also becomes a testing ground to probe higher-order nuclear structure effects and new long- and intermediate-range interactions Berengut et al. (2018); Solaro et al. (2020); Counts et al. (2020); Hur et al. (2022). IS arises mainly due to differences in the masses and nuclear electrostatic potentials of the isotopes, i.e., due to nuclear recoil and field shift effects Yang et al. (2022). The purpose of the atomic electronic structure theory is to calculate these effects and connect experimental observables with fundamental nuclear properties. A very accurate description of IS effects can be achieved in highly charged ions, where electron correlation effects are suppressed and methods of rigorous quantum electrodynamics (QED) are successfully applied (see, e.g., Refs. Zubova et al. (2016); Adkins et al. (2007); Blundell (1993); Anisimova et al. (2022) and references therein). For neutral many-electron atoms, the electronic structure problem becomes very challenging as electron correlation effects are not suppressed and significantly contribute to the atomic factors that describe the IS effects. As a result, approximate treatment of electron correlations usually provides the dominant source of uncertainty in the extracted ms charge radii, negating the experimental achievements. Electron correlation effects also make it difficult to rigorously calculate QED corrections to IS atomic factors, and attempts to treat them in neutral many-electron systems usually result in order-of-magnitude estimates at best.
Until recently Plattner et al. (2023), the value of the ms charge radius for 26mAl was known only from extrapolations Towner and Hardy (2002, 2008). The IS study Plattner et al. (2023) resulted in a significant reexamination of the nuclear ms charge radius of 26mAl. However, the extraction of the radius from these measurements requires input from atomic theory  Heylen et al. (2021); Plattner et al. (2023). The IS between 26mAl and 27Al was measured using collinear laser spectroscopy of the $3s^{2}3p~{}^{2}P_{3/2}\to 3s^{2}4s~{}^{2}S_{1/2}$ transition Plattner et al. (2023), and the uncertainty in the extracted ms charge radii difference between 26mAl and 27Al was strongly dominated by the atomic theory calculations.
In the present paper, we undertake the accurate determination of atomic factors, which allows for a significant reduction in the theoretical and, hence, total uncertainty of the ms charge radius of 26mAl as well as 28Al, 29Al, 30Al, 31Al and 32Al isotopes. To achieve this, we develop a computational scheme to perform high-precision relativistic electronic structure calculations within the coupled cluster theory with inclusion of up to quadruple excitation amplitudes. Furthermore, we develop an effective approach to incorporate leading-order QED effects for these problems with precision comparable to that previously available only for similar transitions in highly charged ions, treated within the rigorous QED approaches.
@@ II.2 QED contribution to the field shift
Here $\theta(R_{\rm cut}-|\mathbf{r}|)$ is the Heaviside step function, $\eta_{nljm}(\mathbf{r})$ are H-like functions (with principal quantum numbers $n\leq 5$), and $R_{\rm cut}$ is a small radius, which can be varied slightly to study the stability of the results obtained. $X_{kljm,k^{\prime}ljm}$ in Eq. (10) are matrix elements of the SE operator over the $h_{kljm}$ functions and, for the reasons mentioned above, to good accuracy they are the linear combinations of the corresponding matrix elements over the H-like functions. Note, that the SE operator is diagonal in $ljm$. The expressions for both diagonal and off-diagonal matrix elements of SE operator can be derived, e.g., within the two-time Green’s function method Shabaev (2002):
$\langle\eta_{n}|\frac{1}{2}(\Sigma^{\rm SE}(\epsilon_{n})+\Sigma^{\rm SE}(\epsilon_{n^{\prime}}))|\eta_{n^{\prime}}\rangle,$ (12)
where $\Sigma^{\rm SE}(\epsilon_{n})$ is the renormalized SE operator. The model SE operator introduced in Ref. Shabaev et al. (2013) or its variant proposed in Ref. Skripnikov (2021) can be used to calculate the SE contribution to the total electronic energy in various systems Shabaev et al. (2013); Tupitsyn et al. (2016); Pašteka et al. (2017); Yerokhin et al. ; Machado et al. (2018); Si et al. (2018); Müller et al. (2018); Kaygorodov et al. (2019); Zaytsev et al. (2019); Shabaev et al. (2020); Skripnikov (2021); Kaygorodov et al. (2021); Savelyev et al. (2022); Kaygorodov et al. (2022); Savelyev et al. (2023); Skripnikov et al. (2021); Athanasakis-Kaklamanakis et al. (2023); Skripnikov and Barzakh (2024). However, they cannot be directly applied to evaluate the SE contribution to the FS constant (7).
@@
where, for brevity, we replaced the set of indices $kljm$ by one index $p$ (or $q$) for elements of the matrix $||X||$ in Eq. (10). We also explicitly indicated the dependence of these matrix elements on the ms nuclear charge radius, and $D_{p,q}^{\langle r^{2}\rangle}$ are elements of the one-particle density matrix in the same basis of functions $h_{p}$, which corresponds to the electronic wave function $\Psi^{\langle r^{2}\rangle}$. Now we have:
@@
To calculate SE contribution due to the first term of Eq. (15), we have to work out a procedure for computing the derivative of the SE matrix elements with respect to the nuclear ms charge radius. In Refs. Milstein et al. (2003, 2004); Yerokhin (2011); Jentschura (2003); Pachucki (1993); Milstein et al. (2002); Eides and Grotch (1997) the nuclear size dependence of the diagonal matrix elements of the SE operator over the functions of H-like ions has been explored and parameterized as follows for $ns,np_{1/2}$ and $np_{3/2}$ states:
@@
$=\frac{\alpha}{\pi}\Delta E_{\rm NS}(nl\frac{1}{2},R)G_{\rm NSE}(nlj,R),$ (16)
where we have omitted the total angular momentum projection in the index of the function $\eta$, $R=\sqrt{\frac{5}{3}}\sqrt{\langle r^{2}\rangle}$, $\Delta E_{\rm NS}$ is the nuclear size (NS) correction to the Dirac energy for the considered state $\eta_{nlj}$, and $G_{\rm NSE}$ is a slowly varying function of $Z$ and $R$ Yerokhin (2011). Note that we use the $nlj$ notation instead of the $njl$ one employed in Ref. Yerokhin (2011). Let us consider the derivative of the matrix element (II.2):
$\frac{d\Delta E^{\rm SE}(nlj,R)}{d\langle r^{2}\rangle}\bigg{|}_{R_{0}}=\Delta E^{\rm SE}(nlj,R_{0})\left[\frac{\frac{d\Delta E_{\rm NS}(nl\frac{1}{2},R)}{d\langle r^{2}\rangle}\bigg{|}_{R_{0}}}{\Delta E_{\rm NS}(nl\frac{1}{2},R_{0})}+\frac{\frac{dG_{\rm NSE}(nlj,R)}{d\langle r^{2}\rangle}\bigg{|}_{R_{0}}}{G_{\rm NSE}(nlj,R_{0})}\right],$ (17)
@@
where $\gamma=\sqrt{1-(\alpha Z)^{2}}$. Our numerical calculations within the hfd code Tupitsyn et al. (2002); Tupitsyn (2003) confirm this analytical result with high accuracy. The second term in the square brackets in Eq. (17) is independent of $n$ for $ns$ and $np_{3/2}$ states according to the explicit expression for $G_{\rm NSE}(nlj,R)$ given in Ref. Yerokhin (2011). A certain dependence on $n$ does exist for $np_{1/2}$ states due to the dependence of the parameter $a_{21}=-2(n^{2}-1)/n^{2}$, which enters into the expression for $G_{\rm NSE}(nlj,R)$ (see Eq. (11) of Ref. Yerokhin (2011)). To explore this dependence, we set $\sqrt{\langle r^{2}\rangle}\approx(2Z)^{1/3}+0.57$ fm approximately following the empirical dependence of $\sqrt{\langle r^{2}\rangle}$ on the nucleon number  Johnson and Soff (1985). We found that non-negligible dependence occurs in the region of $Z=36-40$. If we are outside this range, one can write:
@@ IV Results and discussion
Total, this work $76.81(12)$ $76.80(6)$ $70.11(13)$ $70.10(8)$
Refs. Filippin et al. (2016); Heylen et al. (2021) $76.45$$(1.95)$a $76.20$$(2.20)$
@@
It is evident that the small uncertainties in the non-QED calculations of the atomic IS factors necessitate the consideration of the QED effects. These effects are typically overlooked in theoretical IS studies of many-electron neutral atoms due to uncertainties arising from the incomplete treatment of the correlation effects. Of particular interest is the SE QED contribution $F^{\rm SE}$ to the FS constant, $F^{\rm SE}$ which was found to be larger than the VP contribution (compare lines “QED-SE” and “QED-VP” of Table 5). In Section II.2, we proposed an effective method for performing such calculations for many-electron systems, which allows the use of the relativistic coupled cluster methods. The challenge arises from the lack of exact knowledge of the contribution of the finite nuclear size effect $\Delta E^{\rm SE}(nlj,R_{0})$ in Eq. (19). While the literature provides calculations for $Z=10$, 15, 20, etc. Shabaev et al. (2013), it does not cover $Z=13$. In the section II.2 we mentioned that the interpolation procedure can be used. The good agreement between the proposed technique and the rigorous QED treatment, as shown in Table 1 for Li-like ions, suggests that interpolation is viable. However, since Al is a light atom with $Z=13$, the interpolation scheme for it is based on only one point with a smaller $Z$ ($Z=10$). To verify the reliability of our approach, three calculations of the field shift constant were performed for the $3s^{2}3p~{}^{2}P_{1/2}\to 3s^{2}4s~{}^{2}S_{1/2}$ transition in Al: (i) employing Eqs. (15), (20) and the matrix elements $X^{\rm NS}_{kljm,k^{\prime}ljm}$ interpolated using the data from Ref. Shabaev et al. (2013), along with the analytical expressions for the derivative coefficients $M(lj,R_{0})$ obtained with the use of Eq. (17); (ii) using Eq. (15) with the directly calculated finite nuclear size contributions to the SE matrix elements and numerically computed $dX_{p,q}^{\langle dr^{2}\rangle}/d\langle r^{2}\rangle$ matrix elements within the ab initio QED methods Yerokhin and Shabaev (1999); Malyshev et al. (2022); Sapirstein and Cheng (2023); (iii) directly applying Eq. (7). In the latter case, two sets of the SE matrix elements over the H-like functions were specially calculated for Al for two different nuclear ms charge radii. Consequently, $F^{\rm SE}$ was calculated directly as a numerical derivative. The approach (iii) is the most accurate, while the approach (i) is the most approximate but relies solely on the QED data published in the seminal paper Shabaev et al. (2013) and does not require any new QED calculations. This significantly simplifies the problem of calculating the QED contribution, since an accurate evaluation of the SE matrix elements is not a trivial issue.